OpenStudy (anonymous):

The sum of lexie's age and Melissa's age is 48. Lexie's age is 6 years less than twice Melissa's age. Find Lexie's and Melissa's ages. a. Lexie is 32 years old, Melissa is 16 years old. b. Lexie is 42 years old, Melissa is 6 years old. c. Lexie is 26 year old, Melissa is 22 years old. d. Lexie is 30 years old, Melissa is 18 years old.

OpenStudy (anonymous):

Whole lot of brains here

OpenStudy (anonymous):

l+m is 48, 2m-6 =l

OpenStudy (anonymous):

d. Lexie is 30 years old, Melissa is 18 years old.

OpenStudy (anonymous):

d.

OpenStudy (anonymous):

d

OpenStudy (anonymous):

let lexie's age be x and melissa's age be y x+y=48 x=2y-6 substitute x in 1st equation 2y-6+y=48 3y=54 y=18 substitute y x=2*18-6 x=30

OpenStudy (anonymous):

thank you so much (: